Light having a wavelength in vacuum of 600 nm enters a liquid of refractive index 2.0. In this liquid, what is the wavelength of the light? (c = 3.0 × 108 m/s)?

a) 1200 nm

b) 150 nm

c) 300nm

d) 200nm

e) 600nm
Answer #1

"lambda = c/v" is not using velocity. Depending on where you saw this, that "v" is the Greek letter nu, which means frequency (some sources use "f").

What you want here is to replace "c" in the second equation with your new velocity, after it has slowed down. Frequency does not change when light enters a new medium, so you'll have:

λ(initial) = c/f = 600 nm
λ(final) = v/f = ?

Divide the two equations and you simply get :
λ(i) / λ(f) = c / v = n

which yields:

λ(f) = λ(i) / n = 300 nm.

Ans- c) 300nm